Atoms are typically spaced about 0.2 nm apart. At that spacing, how many atoms would be found in a line [02] inches long?

(b) How many kg of water are in 5.17 cubic inches (in3)?
Answer #1

d = spacing between two adjacent atoms = 0.2 nm = 0.2 x 10-9 m

D = length of the line = 2 inch = 2 x 0.0254 m = 0.0508 m

n = number of atoms

Number of atoms is given as

n = D/d

n = 0.0508/( 0.2 x 10-9 )

n = 2.54 x 108

b)

V = Volume of water = 5.17 in3 = 5.17 (0.0254)3 m3 = 8.5 x 10-5 m3

\rho = density of water = 1000 kg/m3

mass of water is given as

m = \rho V

m = 1000 x 8.5 x 10-5

m = 0.085 kg